According To A Survey,Only 15% Of Customers Who Visited The Website Of A Major Retail Store Made A Purchase.Random Samples Of Size 50 Are Selected.What Proportion Of The Samples Will Have Less Than 15% Of Customers Who Will Make A Purchase After Visiting The Website?
1
Expert's answer
2021-07-21T17:05:39-0400

"P(p<0.15)"

"P(p<0.15)=P(z<\frac{0.15-0.15}{\sqrt{0.15(1-0.15}{50}})"

"=P(z<0)=0.5"

0.5 or 50% of the Samples Will Have Less Than 15% Of Customers Who Will Make A Purchase After Visiting The Website.